1.      You decided not to hedge BusBoard’s currency exposure. If the expected final sales volume is 2,500, what are your total revenues

a)  if the exchange rate (bid-ask) remains at $0.76/C$ - $0.78/C$? Let’s call this the baseline scenario.

b)  if the investors consider the Canadian dollar a safe haven currency during the pandemic? How does this compare to the baseline case?

c)  if the investors consider the U.S. dollar a safe haven currency during the pandemic? How does this compare to the baseline case?


1
Expert's answer
2021-04-29T10:30:33-0400

1. If the expected final sales volume is 2,500, then your total revenues are:

a) if the exchange rate (bid-ask) remains at $0.76/C$ - $0.78/C$, then TR = 0.76×2,500 = $1,900.

b) if the investors consider the Canadian dollar a safe haven currency during the pandemic, then total revenue will decrease.

c) if the investors consider the U.S. dollar a safe haven currency during the pandemic, then the total revenue will increase.